OpenStudy (anonymous):

(tan+cot)^2 = sec^2 + csc^2 (working from one side only to get the other side)

OpenStudy (anonymous):

is it sec^2+csc^2 or sec^2*csc^2?

OpenStudy (anonymous):

sec^2+csc^2

OpenStudy (anonymous):

(x+1/x)^2=x^2+2+1/x^2 so here tan^2+1+cot^2+1=sec^2+csc^2